OpenStudy (anonymous):

multiply 499.99*.0825 show your work help please

OpenStudy (anonymous):

simply multiply 49999 with 825 and then move the decimal to the left 6 times

OpenStudy (amoodarya):

also 499.99*.0825 =(500 -0.001)*.0825 = 500 *0.0825 - 0.01 *0.0825 =41.25 -0.000825